# High-performance Vector-length Agnostic Quantum Circuit Simulations on ARM Processors

### v1

@@ VI Experimental Setup
We use four ARM platforms (Table II) to assess the portability across NVIDIA Grace, AWS Graviton3, and Fujitsu A64FX processors. Grace CPU is based on ARM Neoverse-V2 and two testbeds are used: a single Grace Hopper node with 576 GB memory (the default Grace platform) and a node on the Jupiter supercomputer [18] that consists of four Grace Hopper superchips interconnected by cache-coherent NVLinks to form a four-NUMA system with a total of 864 GB memory. Graviton CPU is based on ARM Neoverse-V1 and runs in a VM on AWS EC2 instance. A64FX is the world’s first processor that implements SVE and was used on Fugaku Supercomputer [19]. The software environment runs Linux kernel 5.14, OpenMP, Python 3.9.18, and pybind11 (v2.13.6) for Python binding to C++. Quantum circuits are constructed and simulated using Cirq 1.3.0.
@@ Acknowledgment
This research is supported by the European Commission under the Horizon project OpenCUBE (101092984). This work was supported by the Lawrence Livermore National Laboratory LDRD Program under 25-ERD-016.

### v2

@@ VI Experimental Setup
We use four ARM platforms (Table II) to assess the portability across NVIDIA Grace, AWS Graviton3, and Fujitsu A64FX processors. The Grace CPU is based on ARM Neoverse-V2 and two testbeds are used: a single Grace Hopper node with 576 GB memory (the default Grace platform) and a node on the JUPITER supercomputer [23] that consists of four Grace Hopper superchips interconnected by cache-coherent NVLinks to form a four-NUMA system with a total of 864 GB memory. The Graviton CPU is based on ARM Neoverse-V1 and runs in a VM on AWS EC2 instance. A64FX is the world’s first processor that implements SVE and was used on Fugaku Supercomputer [17]. The software environment runs Linux kernel 5.14, OpenMP, Python 3.9.18, and pybind11 (v2.13.6) for Python binding to C++. Quantum circuits are constructed and simulated using Cirq 1.3.0.
@@ Acknowledgment
This research is supported by the European Commission under the Horizon project OpenCUBE (101092984). This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 under LDRD Project 25-ERD-016. LLNL-CONF-2016001. Access to JUPITER was provided through the JUPITER Research and Early Access Program (JUREAP); JUPITER is funded by the EuroHPC Joint Undertaking, the German Federal Ministry of Research, Technology and Space, and the Ministry of Culture and Science of the German state of North-Rhine Westphalia.
Appendix: Artifact Description/Artifact Evaluation
Artifact Description (AD)
X Overview of Contributions and Artifacts
X-A Paper’s Main Contributions
$C_{1}$
$C_{2}$
We propose a VLA design for quantum state-vector simulations and optimization techniques, including VLEN-adaptive memory access, buffering, and fine-grained loop control.
$C_{3}$
We provide an implementation in Google’s Qsim, and evaluate five quantum circuits of up to 36 qubits across ARM processors, achieving up to 4.5 $\times$ speedup on A64FX, 2.5 $\times$ on Nvidia Grace, and 1.5 $\times$ on Graviton and high scalability up to 288 threads on the JUPITER supercomputer node.
$C_{4}$
X-B Computational Artifacts
$A_{1}$
https://zenodo.org/records/18661525
Artifact ID Contributions Related
Supported Paper Elements
$A_{1}$ $C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ Table 4
Figures 2, 6-15
XI Artifact Identification
XI-A Computational Artifact $A_{1}$
Relation To Contributions
The artifact A1 is the main source code of this work. By setting backend choice, it can evaluate the capability of auto-vectorization by the compiler, supporting the contribution C1. It provides the implementation of SVE intrinsics version with VLA-specific optimizations, supporting C2 and C3. The lightweight libperf library is designed to support C4.
Expected Results
The runtimes over five circuits are similar in the scenarios with/without auto-vectorization. Compared to the runtime obtained by auto-vectorization, the optimized time by SVE QSim reaches 1.5-4.5×, depending on different SVE vector lengths. Counting PMU events, the higher proportion of vector instructions and vector activites can be observed.
Expected Reproduction Time (in Minutes)
The longest expected computational time traversing five circuits in 32 qubits per time is around 300 seconds in SVE QSim and 1200 seconds in Navie QSim. The time requires doubles with each additional qubit.
Artifact Setup (incl. Inputs)
Hardware
Our implementation can be performed on ARM CPU equipped with SVE. The memory capability requires more than $2^{(N+3)}$ B for $N$-qubit simulation. The experiments were performed on Nvidia Grace, Graviton and A64FX CPU.
Software
Python 3.9+ with Cirq, Numpy, pybind11, and setuptools package; cmake 3.26.5; GNU compiler (GCC12 and above) is used for experiments in this paper. To enable perf event, perf v5.14.0 is needed and perf_event_paranoid value of less than 1.
Installation and Deployment
Build the project by pip install -e .
Artifact Execution
To obtain the comparison results of non/auto-vectorized/, set CMAKE_CXX_FLAGS to one of: NonVec, SVE, Neon options; Recompile and navigate to benchmark directory, executing INSTR=4 run_qsim_leq32.py with the experimental parameter 31 qubits, fused 3. For SVE QSim performance, INSTR=3 switches to SVE intrinsics backend ./benchmark/run_multibench.sh specifies the the experimental parameters running on Grace CPU and similar execution steps for other ARM CPUs when adequate memory and threads. run_qsim_GPU.py can execute quantum simulation with CUDA.
Artifact Analysis (incl. Outputs)
Execution time and PMU event counts will be printed out and can be saved to a log file. ./benchmark/parse_output.sh *.log can extract PMU event counts and calculate the average. The results in the paper represent the average of five runs.
Artifact Evaluation (AE)
XI-A Computational Artifact $A_{1}$
Artifact Setup (incl. Inputs)
Please download and uncompress the source code,
1wget "https://zenodo.org/records/18661525/files/qsim-arm-sve-intrinsics.zip?download=1" -O qsim-arm-sve-intrinsics.zip
2unzip qsim-arm-sve-intrinsics.zip
3cd qsim-arm-sve-intrinsics
Please create the python environment and install qsim and its dependency
1python -m venv qsim
2source qsim/bin/activate
3cd libperf/ && make && cd ..
4pip install -e .
Change ./pybind_interface/basic/CMakeLists.txt Line 7 CMAKE_CXX_FLAGS to non/auto-vectorized compiler flags,
  • Non-vectorized: -O3 -fno-tree-vectorize -fno-tree-loop-vectorize -fno-tree-slp-vectorize
  • Auto-vectorized(default): -march=armv8.5-a+sve -mcpu=neoverse-v2
Artifact Execution
The format to execute the simulator,
1INSTR=<3/4> python benchmark/<script>.py -b <circuit> -q <qubits> -t <threads> -f <2/3/4/5> -r <iterations>
The usage of provided scripts,
  • run_qsim_leq32 is running a circuit with less than or equal to 32 qubits.
  • run_qsim_gt32.py is running a circuit with greater than 32 qubits.
  • correctness.py is validation of the results.
  • run_qsim_GPU.py is running a circuit in signle GPU.
INSTR=3 runs QSIM SVE; INSTR=4 runs the naive version of Qsim by non/auto-vectorized compiling. -b <circuit> has the option: grover, ghz, QRC, qft, qv.